### v1

# RainPro-8: An Efficient Deep Learning Model to Estimate Rainfall Probabilities Over 8 Hours

Abstract
We present a deep learning model for high-resolution probabilistic precipitation forecasting over an 8-hour horizon in Europe, overcoming the limitations of radar-only deep learning models with short forecast lead times. Our model efficiently integrates multiple data sources - including radar, satellite, and physics-based numerical weather prediction (NWP) - while capturing long-range interactions, resulting in accurate forecasts with robust uncertainty quantification through consistent probabilistic maps. Featuring a compact architecture, it enables more efficient training and faster inference than existing models. Extensive experiments demonstrate that our model surpasses current operational NWP systems, extrapolation-based methods, and deep-learning nowcasting models, setting a new standard for high-resolution precipitation forecasting in Europe, ensuring a balance between accuracy, interpretability, and computational efficiency.
1 Introduction
Recent advances in artificial intelligence, especially in computer vision and natural language processing, have generated significant interest in deep learning for weather forecasting (Rasp et al., 2024; An et al., 2024). Deep learning excels at handling complex, large-scale, high-dimensional datasets, making it well-suited for capturing intricate, nonlinear patterns in spatio-temporal systems (Manzhu Yu and Li, 2024). Although deep learning has achieved remarkable success in both nowcasting (Gao et al., 2024b; Gong et al., 2024) and medium-range forecasting (Lam et al., 2023; Price et al., 2024), significant challenges remain. Deep learning models for nowcasting are often limited to very short lead times (up to two hours). In contrast, medium-range models, which predict broader atmospheric dynamics for up to 10 days, typically operate at coarser resolutions and are influenced by precipitation-specific biases in the training datasets (Lavers et al., 2022). As a result, they struggle to capture small-scale precipitation features, like local showers, often leading to the exclusion of precipitation forecasts in medium-range models (Lam et al., 2023).
This work addresses the challenge of forecasting precipitation for up to 8 hours at high spatio-temporal resolutions of 2 km per pixel and 10-minute intervals, bridging the gap between nowcasting and medium-range forecasting. Forecasting over an 8-hour horizon is critical for timely predictions that help mitigate risks like flooding and optimize resource management in sectors such as agriculture, energy, and transportation. This task is particularly difficult due to the stochastic and sparse nature of precipitation, especially over extended lead times. Designing a deep learning model for this task requires addressing skewed and intermittent precipitation distributions, incorporating multi-sensor data, and ensuring robust uncertainty quantification.
Most existing deep learning-based nowcasting methods focus on generating radar-like precipitation forecasts (Ravuri et al., 2021; Gao et al., 2024b), but these become increasingly challenging and less accurate with longer lead times. Instead, probabilistic models emphasize uncertainty estimation and provide a more accurate and practical representation of rainfall at varying intensities. The MetNet family of models (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023) represents state-of-the-art deep learning-based systems for probabilistic precipitation forecasting. These models produce high-resolution forecasts for 8-24 hours in the United States, outperforming operational NWP systems. They achieve efficient probabilistic forecasting using cross-entropy loss on precipitation bins, requiring only one forward pass rather than an ensemble. However, this approach neglects the ordinality between bins. In addition, they rely on lead-time conditioning, where the model generates forecasts for only one lead time at a time. While this approach is fully parallelizable across multiple GPUs, it significantly increases inference computational demands. Moreover, training MetNet requires substantial computational resources and is limited to the United States.
We propose RainPro-8, an efficient deep learning model to estimate rainfall probabilities over 8 hours in Europe by determining the probability of different levels of precipitation at a given location and time. RainPro-8 is an efficient model based on MetNet-3 (Andrychowicz et al., 2023) that uses less than 20% of MetNet-3’s training parameters to achieve 8-hour high-resolution precipitation forecasting across Europe — a region characterized by diverse climates, complex terrain, and highly variable precipitation dynamics (Ehmele et al., 2020). We also redesign the training objective to explicitly account for the ordinality between precipitation bins. Lastly, we demonstrate that it is possible to generate all lead times simultaneously by downweighing later lead times during training, greatly enhancing inference efficiency and improving temporal consistency. Our model demonstrates substantial improvements over traditional precipitation forecasting methods, including NWP models and extrapolation-based techniques, as well as deep-learning nowcasting models, across various rain intensities for the next 8 hours.
Our main contributions are as follows:
  • We introduce an efficient model architecture and training strategy that learns from multi-source data with varying temporal and spatial resolutions, accounts for the ordinality of precipitation classes, and outputs all lead times simultaneously.
  • We evaluate our approach on a European dataset we created specifically for 8-hour precipitation nowcasting and demonstrate that our method outperforms existing operational methods by 65%, as well as deep-learning nowcasting models.
  • We perform ablation studies and model attribution analysis to assess RainPro-8 performance and the effects of key design choices and input features.
2 Related Work
Traditional methods for weather forecasting are mostly numerical weather prediction (NWP) models, which simulate atmospheric dynamics using mathematical equations, also as ensembles over multiple simulations (Toth and Kalnay, 1997). NWP models demand significant compute, especially for ensembles, which restricts their spatial and temporal resolution. Their spin-up time results in poor performance for short lead times (Ma et al., 2021). Extrapolation-based methods, e.g. PySTEPS (Pulkkinen et al., 2019), RainyMotion (Ayzel et al., 2019) are limited to reduced forecast lengths due to their assumptions on constant motion and intensity (van Nooten et al., 2023).
Deep learning formulates precipitation nowcasting as a spatio-temporal prediction (An et al., 2024). Approaches in this area include the use of convolutions in recurrent networks (Shi et al., 2015, 2017; Wang et al., 2023b; Ma et al., 2024), the U-Net architecture (Ayzel et al., 2020; Fernández and Mehrkanoon, 2021; Zhang et al., 2023; Trebing et al., 2021), transformers (Gao et al., 2024b; Wu et al., 2024), diffusion models (Leinonen et al., 2023; Gao et al., 2024a; Gong et al., 2024), and adversarial training (Ravuri et al., 2021). Additionally, some research focuses on improving the loss function to tackle issues such as data imbalance, blurriness, and sparsity in precipitation nowcasting models (Xu et al., 2024a; Ko et al., 2022; Yan et al., 2024). Despite extensive research in this area, most efforts have focused on short-term forecasts (1 to 3 hours) using only radar, where the reduced lead time eliminates the need for multiple data sources, large contexts, or accounting for the higher uncertainty associated with 8-hour predictions.
Limited work exists in precipitation forecasting for lead times beyond 3 hours. The Weather4Cast competition (Gruca et al., 2022) tackles 8-hour precipitation forecasting from satellite alone, but it has shown limited success compared to traditional methods (Li et al., 2023) due to challenges in estimating current precipitation from satellite imagery and overcoming resolution issues compared to radar. For 6-hour precipitation nowcasting, NPM (Park et al., 2024) relies solely on satellite data without data fusion, while Nowcast-to-Forecast (An, 2023) combines satellite and radar but is limited by a biased rainy-day dataset. Kim et al. (2024) only compare hourly forecasts to extrapolation baselines, which are ineffective for longer-term predictions. TAFFNet (Wang et al., 2023a) provides 12-hour predictions but neglects uncertainty and relies on radar and NWP data at equal resolutions, which are not always available.
MetNet-3 (Andrychowicz et al., 2023) is a leading deep learning model for precipitation forecasting, delivering 24-hour forecasts with high spatio-temporal resolution in the U.S. by integrating radar, weather station, satellite imagery, assimilated weather states, and other data in a transformer-based architecture. However, its training requires significant time and resources, involving hundreds of Tensor Processing Units (TPUs) for multiple days. Its reliance on high-quality data available only for the contiguous U.S. and the lack of public access to its code and data restrict its broader use. In this work, we show how to adapt its approach to problem formulation, data preprocessing, and training optimization for smaller and more efficient models and specific data requirements.
Another line of work focuses on medium-range forecasting. Medium-range models like GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), or GenCast (Price et al., 2024) rely on reanalysis datasets like ERA5 (Hersbach et al., 2020), which suffer from low resolution, delays, and biases, especially in surface variables and precipitation (Lavers et al., 2022). Precipitation, highly variable and challenging to predict, is underrepresented in such models (Rasp et al., 2024), with most failing to integrate critical observational data like radar. While some models, like NeuralGCM (Yuval et al., 2024), incorporate global precipitation predictions, they lack the high resolution and operational frequency needed for our setting. Other work (Xu et al., 2024b) has tried to address the gap between nowcasting and medium-range forecasting, but model performance drops at high temporal resolutions and focuses only on low precipitation.
3 Method
3.1 Probabilistic Precipitation Forecasting
Precipitation forecasting is a spatiotemporal prediction problem, where the goal is to predict $T_{\text{out}}$ future radar frames $Y$ based on a sequence of $T_{\text{in}}$ past radar frames $X$:
$X=[R_{t}]^{0}_{t=-T_{\text{in}}+1}\in\mathbb{R}^{T_{\text{in}}\times H\times W}\qquad Y=[R_{t}]^{T_{\text{out}}}_{t=1}\in\mathbb{R}^{T_{\text{out}}\times H\times W},$ (1)
where $R_{t}$ represents the precipitation intensity at timestep $t$ based on radar maps of size $H\times W$.
Radar data alone is not enough for accurate precipitation forecasting over long periods (e.g., up to 8 hours) because its reliance on ground-based systems limits its coverage and it cannot capture atmospheric conditions beyond water vapor. To enhance forecasting accuracy for longer lead times, additional data sources like satellite or NWP provide broader coverage and a more comprehensive representation of the atmosphere. Such additional sources introduce challenges due to differences in temporal frequency and spatial resolution, requiring careful preprocessing and alignment as detailed in Appendix A. The generalized input $X$, with heterogenous sources Sources, is given by:
$X=\bigcup_{S\in Sources}[S_{t}]^{0}_{t=-T_{\text{in\_s}}+1},$ (2)
where $S_{t}\in\mathbb{R}^{C_{s}\times H_{s}\times W_{s}}$ represents a frame from the data source $S$, characterized by a specific number of channels, spatial dimensions, and resolution.
In addition to spatiotemporal accuracy, another critical priority is generating probabilistic forecasts that address the question: What is the probability of a specific amount of rainfall at a given location and time? Commonly, quantifying uncertainty involves generating multiple forecasts for ensembles to calculate probabilities, but this approach demands substantial computational resources. Instead of relying on ensemble methods, our approach directly predicts the probability distribution of precipitation intensities, similar to MetNet 1-3 (Sønderby et al., 2020; Espeholt et al., 2022; Andrychowicz et al., 2023). Our goal is to generate accurate probability maps to better capture uncertainty and variability in precipitation patterns, instead of radar-like outputs, which become less reliable over longer lead times. To model probabilities for different precipitation intensity classes $I$, we redefine the target as probability maps for each precipitation intensity class:
$Y=\bigcup_{t=1}^{T_{\text{out}}}\bigcup_{c=1}^{|I|}P_{t,c}\in\mathbb{R}^{T_{\text{out}}\times|I|\times H\times W},$ (3)
where $I$ is the set of intensity classes that divides the possible precipitation intensities into ranges or bins, and $P_{t,c}$ is the probability map for $R_{t}$ with respect to class $I_{c}$.
In the case of MetNet, $P_{t,c}=P(R_{t}\in I_{c})$, which predicts the probability of precipitation intensity falling within specific predefined ranges. Given this definition, MetNet is trained with the cross entropy loss. However, this approach ignores the intrinsic order of the intensity classes, which we aim to solve with the proposed Ordinal Consistent loss.
3.2 Ordinal Consistent Loss
To generate probabilistic forecasts that preserve ordinality among precipitation classes, we model $P_{t,c}$ as $P(R_{t}\geq\min(I_{c}))$. To ensure monotonicity ($P_{t,c}\leq P_{t,c-1}$), we reformulate $P_{t,c}$ using Bayes’ theorem, given that $P(R_{t}\geq\min(I_{c-1})|R_{t}\geq\min(I_{c}))=1$, and redefine the model outputs:
$P_{t,c}$ $=P(R_{t}\geq\min(I_{c}))=\frac{P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))\times P(R_{t}\geq\min(I_{c-1}))}{P(R_{t}\geq\min(I_{c-1})|R_{t}\geq\min(I_{c}))}$ (4)
$=P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))\times P_{t,c-1}$
By having the model output $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ within the range $[0,1]$ for each intensity class $c$, we guarantee monotonic probabilities $P_{t,c}$ that respect the inherent order of precipitation classes. For the lowest intensity class, the model outputs $P(R_{t}\geq\min(I_{1})$. This approach, as proposed by Fernandes and Cardoso (2018) for semantic segmentation, improves both interpretability and consistency in the generated forecasts.
For any given class, the predicted probability $P_{t,c}$ is computed as the cumulative product of the model’s outputs for all preceding classes:
$P_{t,c}$ $=P(R_{t}\geq\min(I_{c})\mid R_{t}\geq\min(I_{c-1}))\times P_{t,c-1}$ (5)
$=\prod_{j=2}^{c}P(R_{t}\geq\min(I_{j})\mid R_{t}\geq\min(I_{j-1}))\times P(R_{t}\geq\min(I_{1}))$
The loss uses target binary masks indicating whether $R_{t}\geq\min(I_{c})$ for each class and timestep. Binary Cross Entropy (BCE) is used to compute the loss between the target masks and predicted probabilities. This calculation is performed without reduction to yield a per-pixel loss for each spatial location, class, and timestep.
To enhance training, the ordinal consistent mask (OC mask), computed for every sample as $R_{t}\geq\min(I_{c-1}))$, is incorporated. It ensures that for each intensity class $c$, the loss is averaged only over pixels where the previous class is activated, helping the model take advantage of the implicit ordinality. The loss is also not averaged over prediction pixels that lack ground truth values, typically missing due to the limited coverage of radar data. The loss function, based on BCE with the OC mask, is defined as:
$\mathcal{L}$ $=\text{mean}\left(\text{loss}(t,c,h,w)\right)\quad\text{if}\quad\text{mask}(t,c,h,w)\neq 1,$ (6)
[…]
$\quad\quad\quad\quad\text{mask}(t,c,h,w)=\mathds{1}\left(R_{t}(h,w)=-1\right)\lor\mathds{1}\left(R_{t}(h,w)<\min(I_{c-1})\right),$
where $R_{t}(h,w)=-1$ in areas without radar coverage.
3.3 Single-Pass Predictions
Precipitation nowcasting becomes significantly harder as lead times increase, which can negatively impact the model training. Auto-regressive models focus on one step at a time for stability but struggle when input sources differ from targets. MetNet tackles this with lead time conditioning, which supports multiple sources, but this approach significantly reduces inference efficiency and may cause temporal inconsistencies between lead times.
Instead, we make single-pass predictions to generate all forecast timesteps in one forward pass, reducing resource requirements, accelerating training convergence and inference, and improving temporal consistency in the outputs. The model encodes timestamps into the channel dimension, with the output $B(TC)HW$ reshaped to $B\,T\,C\,H\,W$, where $T$ represents all timesteps.
Inspired by MetNet-3’s lead-time sampling strategy, we introduce lead time weights. With lead-time sampling, lead times are sampled using a biased scheme based on an exponential distribution, which results in more frequent use of samples with shorter lead times during training, giving them a larger influence on model loss. In contrast, our approach includes all lead times in every training sample, so we incorporate lead time-dependent weights into the loss function to reduce the weighting of longer lead times. This ensures the model still prioritizes earlier lead times during training, improving performance across the entire range of lead times, including the more challenging longer ones. Therefore, the pixel-wise loss function is multiplied by lead-time weight $W_{t}$, normalized from an exponential distribution with a given decay rate $\alpha$, determining the relative weight assigned to the first timestep compared to the last one.
$\textrm{weights}_{\textrm{exp}}[t]=\exp(-\alpha\times t)\quad\textrm{weights}_{\textrm{norm}}[t]=\frac{\textrm{weights}_{\textrm{exp}}[t]}{\sum\textrm{weights}_{\textrm{exp}}[t]}\quad W_{t}=\frac{\textrm{weights}_{\textrm{norm}}[t]}{\overline{\textrm{weights}_{\textrm{norm}}}}\$ (7)
where $\textrm{weights}_{\textrm{exp}}$ is the weights based on exponential decay, $\textrm{weights}_{\textrm{norm}}$ is the normalized weights, and $W_{t}$ is the rescaled weights so the total loss scale remains consistent with the unweighted case.
3.4 Architecture
We design our architecture to forecast precipitation for a $512\times 512\textrm{km}^{2}$ patch. Assuming an average precipitation displacement rate of 1km per minute (Sønderby et al., 2020), we use 512km of spatial context on every side of the target patch ($\textrm{1536}^{2}\textrm{km}^{2}$) to provide our model with enough input context for all 8-hours of lead time.
Inspired by MetNet-3 (Andrychowicz et al., 2023), we use a U-Net (Ronneberger et al., 2015) with MaxViT (Tu et al., 2022) blocks to efficiently process multi-resolution data (Figure 1). Key differences include single-pass predictions without lead time conditioning (Section 3.3), early downsampling in the encoder, halving the internal channels, and removing topographical embeddings, all contributing to a reduced parameter count of 36.7M from the original 227M.
Figure 1: RainPro-8, optimized for reduced parameter count and improved efficiency, enabling simultaneous output of all timesteps and integration of data sources with varying resolutions.
The architecture begins with data fusion on the encoder, using Space-to-Depth convolutions (Sunkara and Luo, 2023) and ResNet blocks (He et al., 2016). Input data is transformed from $BTCHW$ to $B(TC)HW$ and merged with sources at matching resolutions. To balance efficiency and performance, we only use the full input context of 512km at 8km and 16km inputs, and use a lower input context of 256km at 4km resolution. The encoder aggregates features for each resolution into a low-resolution representation while preserving skip connections for the U-Net decoder. We apply downsampling before ResNet blocks instead of after to improve memory efficiency.
Next, the low-resolution representation is processed by 12 MaxViT blocks. MaxViT blocks, based on Vision Transformer (Dosovitskiy et al., 2021), combine local neighborhood and global gridded attention aggregate information across the full $\textrm{1536}^{2}\textrm{km}^{2}$ input patch. The decoder, using Transposed Convolutions (Long et al., 2015) and ResNet blocks, reconstructs high-resolution probability maps from the MaxViT output, while leveraging encoder skip connections. All forecast timesteps and intensity classes are generated by outputting $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ (Section 3.2) in a single forward pass with timestamps encoded into the channel dimension (Section 3.3).
To ensure geographical alignment, representations are padded and cropped at each resolution stage. For example, 4km representations are padded equally on all sides before concatenating with 8km inputs, and the MaxViT output is cropped to the target region with extra context for upsampling, later cropped to $\textrm{512}^{2}\textrm{km}^{2}$ for the final output.
4 Experiments
RainPro-8 is implemented with PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon and team, 2025), trained for 100k steps using a batch size of 16, validating every 2,000 steps. Training uses a static learning rate of 3e-4, AdamW optimizer with a weight decay of 0.1 and betas $(0.9,0.999)$, Exponential Moving Average (EMA) decay of 0.99975, dropout of 0.1, maximum stochastic depth (Huang et al., 2016) of 0.2, and lead-time decay rate of 10. We use 256 channels throughout the entire network, totaling 36.7 million parameters. Any other network hyperparameters follow those of MetNet-3 (Andrychowicz et al., 2023). The model with the lowest validation loss is selected. Training is performed on an NVIDIA H100 80GB SXM5 GPU and takes approx. 13 hours to complete.
We use multiple data sources as input to our precipitation forecasting model, each matched to its supported spatial resolutions. RainViewer [^1] https://www.rainviewer.com/ radar composites serve as high-resolution ground truth at 4km and 8km resolutions to capture local detail and broader context. EUMETSAT [^2] https://user.eumetsat.int/data/satellites/meteosat-second-generation satellite imagery provides cloud-related features at 8km resolution, and NOAA’s GFS [^3] https://registry.opendata.aws/noaa-gfs-bdp-pds provides atmospheric variables and precipitation forecasts at 16km resolution. Topographical information from Copernicus DEM [^4] https://registry.opendata.aws/copernicus-dem is incorporated at 4km resolution. All data are resampled for spatial and temporal alignment and undergo normalization, clipping, and binning to handle varying scales and precipitation skewness. The data cover one year and over one million samples, with defined training, validation, and testing splits to ensure reliable model evaluation. Details on data, preprocessing, and samples in App. A.
We use Critical Success Index (CSI) (Schaefer, 1990) at different thresholds, a standard accuracy metric in precipitation nowcasting (Andrychowicz et al., 2023; Gao et al., 2024b; Gong et al., 2024), and Continuous Ranked Probability Score (CRPS) (Hersbach, 2000), which assesses alignment of predicted probability distribution with observed values, rewarding sharp and reliable predictions. Fractions Skill Score (FSS) (Roberts and Lean, 2008) accounts for intensity shifts and offers tolerance to translation and deformation. Frequency Bias Index (FBI) (Termonia et al., 2018) quantifies over- and underforecasting, but not forecast quality. We also report Mean Absolute Error (MAE) and Mean Squared Error (MSE), but note that they are sensitive to high frequency of no-rain cases and thus less suitable for skewed distributions (Andrychowicz et al., 2023) (cf. App. B.1). Taking inspiration from the thresholding approach in MetNet (Espeholt et al., 2022), intensity-based metrics use the mean of the highest activated bucket of the cumulative output distribution. A bucket is activated if its predicted probability mass exceeds the corresponding threshold, where thresholds are computed for each bucket and lead time using the validation set. This approach effectively captures rare high-precipitation events, which tend to have lower predicted probability masses.
4.1 Performance Evaluation
We compare our model with global NWP GFS [^5] https://registry.opendata.aws/noaa-gfs-bdp-pds and extrapolation-based PySTEPS (Pulkkinen et al., 2019). PySTEPS extrapolates radar echoes for a few hours with a context of 512km. NWP forecasts are bilinearly interpolated to a 2 km/px resolution and temporally aligned to 10-minute intervals for radar comparison. Direct comparison with deep-learning nowcasting models is challenging due to fundamental limitations in their architectures and input assumptions. Specifically, models such as NowcastNet (Zhang et al., 2023), Earthformer (Gao et al., 2024b), and SimVP (Gao et al., 2022) operate only on radar data with fixed resolution and spatial coverage, and integrating additional data sources would require substantial architectural changes. To ensure a fair comparison, we provide all models with the same input region, including additional context, and evaluate them solely on the target central region. We train these models using their original architectures with MSE loss. SimVP, which forecasts sequences of the same length as its input, requires an autoregressive approach. Inputs are interpolated bilinearly from 4km to 2km for NowcastNet and SimVP. Earthformer, however, requires training at 4km resolution due to attention mechanism constraints, with outputs upsampled to 2km for evaluation. MetNet-3 (Andrychowicz et al., 2023) evaluation presents significant challenges due to its private code, reliance on US-specific data, and substantial computational requirements (227M parameters trained on 512 TPU v3 cores over 7 days). MetNet-3* is our faithful reimplementation of the architecture and training described in the original paper, adapted to our data and compute constraints, presenting a reproducible competitor under fair conditions. It incorporates lead-time conditioning and cross-entropy loss, and it only computes the loss on the high-resolution precipitation maps, not on accumulated rain or NWP initial state since we do not perform densification from weather stations.
As shown in Table 1, RainPro-8 outperforms all competitors on precipitation metrics — Critical Success Index (CSI) and Fractions Skill Score (FSS) — demonstrating strong skill and practical utility in rain forecasting. Frequency Bias Index (FBI) indicates that RainPro-8 slightly overpredicts precipitation, while PySTEPS is least biased. Although NowcastNet, SimVP, and Earthformer achieve lowest MAE and MSE due to their MSE loss in short-term nowcasting, these do not translate into higher CSI and FSS, likely because of domination of the no-rain class in the highly imbalanced dataset. RainPro-8 of performs slightly better than MetNet-3*, and offers additional advantages: a 48×48$\times$48 × inference speedup through single-step prediction and more coherent probability maps, enabled by the ordinal-consistent loss that accounts for class ordinality in probabilistic forecasts.
Table 1: Model comparison over all lead times and intensity thresholds.
Model Radar-only? CSI ($\uparrow$) FSS ($\uparrow$) FBI ($\approx 1$) MAE ($\downarrow$) MSE ($\downarrow$)
ours $\times$ 0.279 0.537 1.262 0.126 1.503
metnet-3* $\times$ 0.270 0.517 1.318 0.132 1.620
gfs $\times$ 0.110 0.253 0.780 0.164 1.453
pysteps $\surd$ 0.149 0.364 0.983 0.162 2.324
nowcastnet $\surd$ 0.108 0.255 0.153 0.114 1.354
earthformer $\surd$ 0.111 0.267 0.163 0.110 1.358
simvp $\surd$ 0.122 0.287 0.189 0.118 1.340
Figure 2 illustrates model performance (CSI) at different rain intensities and lead times up to eight hours (further plots and tables in App. B.2). Our model demonstrates superior performance across lead times and for all rainfall thresholds, ranging from light to heavy rain. The skill gap with extrapolation-based PySTEPS grows as lead times increase due to its assumption of constant motion and intensity. NowcastNet, Earthformer, and SimVP skill gap also grows with lead times due to the lack of additional data sources and probabilistic outputs. In contrast, GFS shows a larger skill gap at shorter lead times because of its longer convergence time. Performance is slightly lower for all models at higher precipitation intensities, likely due to the increased complexity in forecasting extreme weather events, which often involve more chaotic and unpredictable patterns.
Figure 2: Critical Success Index (CSI) across different thresholds and lead times.
Figure 3 shows forecast for a $1000^{2}\textit{km}^{2}$ cropped region of at least 4 target patches of $512^{2}\textit{km}^{2}$, at three lead times (Europe-wide predictions visualized in App. D). Our model effectively captures most areas with rainfall, both in terms of intensity and location, with some blurriness as lead time increases due to growing uncertainty. GFS shows major deviations from the ground truth, with large differences in both intensity and rain-covered areas. PySTEPS, which can only extrapolate the latest radar image, struggles to capture changes in intensity or account for non-linear motion patterns at later lead times. NowcastNet, Earthformer, and SimVP are limited by inability to use more data sources—crucial for reliable 8-hour forecasts—and by MSE-loss, which favors frequent no-rain.
Figure 3: Sample ground truth and forecasts for different models at selected lead times with origin at 2024-01-23 11:20 UTC in cropped region. Dark grey areas indicate regions beyond radar coverage.
4.2 Ablation Studies
Table 2 shows ablation studies on the use of the ordinal-consistent loss function and single-pass predictions, compared to MetNet-3*, which uses non-ordinal cross-entropy loss and lead time conditioning (Andrychowicz et al., 2023). We also study the impact of two design choices in our loss function: class masking and lead time weighting, as well as the contribution of satellite data, topography, and GFS forecasts input. The results confirm the effectiveness of our approach, highlighting the advantages of the ordinal consistent loss with masking and single-pass predictions for generating 8-hour forecasts with lead time weights and all data sources. The slight gains of our model over the cross-entropy-based model are due to the latter enforcing ordinality only at inference time via cumulative probability computation, rather than learning it during training. Compared to models using lead-time conditioning, our single-pass prediction approach not only improves inference speed but also delivers superior performance across metrics. Excluding lead-time weights leads to a noticeable drop in performance, highlighting their importance in the loss. Similarly, the model without OC mask underperforms, likely because the model needs to focus on lower intensity classes before predicting higher ones, as identifying the first change in class activation (indicating expected rain) is more critical. Finally, utilizing all available data is key to achieving top forecasting accuracy.
Table 2: Ablation loss function, timesteps per forward pass, mask, lead time weights, data sources.
Model $\mathcal{L}$ Tsize OCmask OCltw Srcs. CRPS ()($\downarrow$)( ↓ ) CSI ()($\uparrow$)( ↑ ) FSS ()($\uparrow$)( ↑ )
RainPro-8 OC 48 $\surd$ $\surd$ $\surd$ 0.06096 0.2791 0.5367
MetNet-3* CE 1 - - $\surd$ 0.06199 0.2697 0.5173
RainPro-8_ce CE 48 - - $\surd$ 0.06098 0.2787 0.5357
RainPro-8_clt OC 1 $\surd$ - $\surd$ 0.06203 0.2695 0.5191
RainPro-8_noltw OC 48 $\surd$ $\times$ $\surd$ 0.06156 0.2733 0.5258
RainPro-8_nomask OC 48 $\times$ $\surd$ $\surd$ 0.06971 0.2565 0.5151
RainPro-8_sources OC 48 $\surd$ $\surd$ $\times$ 0.06331 0.2538 0.4910
4.3 Probability Maps
Our RainPro-8 model not only outperforms traditional operational methods in generating precipitation intensity maps, it also offers the advantage of quantifying uncertainty in predictions for each rainfall intensity. Figure 4 displays probability maps alongside ground truth for two different rain intensities and lead times, using the same sample and cropped region as in Figure 3, where intensity values were derived from these probability maps. The results show higher uncertainty for longer lead times and lower precipitation intensities, reflecting the inherent challenge of predicting scattered light rain far into the future at high resolution. For higher intensities, the model tends to predict low probabilities due to their low likelihood, particularly at extended lead times. Our model ensures consistency by guaranteeing that the probability of experiencing at least a certain precipitation level is always greater than or equal to the probability of experiencing higher precipitation levels.
Figure 4: Ground truth and RainPro-8 probability map for different rain intensities and lead times, origin at 2024-01-23 11:20 UTC. Dark grey areas indicate regions beyond radar coverage.
4.4 Attribution
We study RainPro-8 using Integrated Gradients (IG) (Sundararajan et al., 2017), as done for MetNet-2 (Espeholt et al., 2022). IG attributes predictions to input features, offering insights into how the model utilizes each data source. Key IG findings (cf. App. C) are: high-resolution recent radar data is crucial for short lead times, while low-resolution radar gains importance around the 4-hour mark due to its broader coverage. Satellite data becomes important for lead times beyond 4 hours, where radar data is less effective. Visible channels have minimal impact, likely because they are restricted to daytime. GFS variables become more influential for longer lead times, with key contributors including wind components, vertical velocity, storm motion parameters, specific humidity, and surface metrics such as pressure and the Best Lifted Index (a measure of atmospheric instability), which are particularly relevant for forecasting cloud movement, precipitation, and thunderstorms. GFS forecasts become increasingly impactful at longer lead times, while less relevant for the first 4 hours.
5 Conclusion & Future Work
We introduce RainPro-8, a deep learning model for precipitation nowcasting in Europe that outperforms existing operational systems and deep learning nowcasting models. RainPro-8 addresses the challenges of forecasting sparse and stochastic precipitation by probabilistic predictions at different precipitation levels. It leverages multiple data sources with an efficient model that forecasts all lead times simultaneously, while considering the ordinality of precipitation levels.
Going forward, RainPro-8 could be extended to be more robust to structural changes in input data and to missing data — the latter is of particular importance in operational settings where some input source could become absent. RainPro-8 uses GFS forecasts as part of its inputs, where it would be interesting to identify alternative data sources, and study automatically learned thresholds.
Appendix A Dataset
A.1 Data Sources
We use RainViewer [^6] https://www.rainviewer.com/ radar composite data over Europe, which consists of high-resolution rainfall intensity measurements at 10-minute intervals and a resolution of 2 kilometers per pixel. Geostationary satellite data is obtained from the European Organisation for the Exploitation of Meteorological Satellites (EUMETSAT) [^7] https://user.eumetsat.int/data/satellites/meteosat-second-generation, capturing cloud-related characteristics that significantly influence precipitation, despite its lower resolution (3-11km) and the challenges of correlating satellite and radar data (Stock et al., 2024). Topographical data is retrieved from the Copernicus Digital Elevation Model (DEM) [^8] https://registry.opendata.aws/copernicus-dem. Atmospheric observations and derived physical parameters are sourced from the Global Forecast System (GFS) [^9] https://registry.opendata.aws/noaa-gfs-bdp-pds, managed by the National Oceanic and Atmospheric Administration (NOAA) in the United States, with a spatial resolution of  28km and update frequency of 6 hours. In contrast, MetNet (Andrychowicz et al., 2023) uses HRRR, with 3km resolution, hourly updates, and US-only coverage.
A.2 Data Preprocessing
Each data source is modified to match the desired resolution for model input, either through upsampling, such as GFS data from 28km/px to 16km/px, or downsampling, like radar data from 2km/px to both 4km/px and 8km/px. The time dimension also presents challenges to consider. Satellite data, for instance, has a 1-hour operational delay, making the most recent timestep available at -60 minutes. For the source gfs_16km, we select 122 channels representing multiple weather variables at different pressure levels (Appendix F). As GFS is only initialized 4 times a day, we use the latest forecast lead time as our initial state. The subsequent lead times form the gfs_forecast_16km, from which we use only the precipitation rate variable. Table 3 summarizes all data sources with the corresponding spatial and temporal details.
Table 3: Outputs and inputs for the precipitation forecasting model.
Variable Source
Size
Res.
(km/px)
Size
Timesteps
Channels
target_2km RainViewer 256 2 512 [10,20,…,480] 1
radar_4km RainViewer 256 4 1024 [-60,-50,…,0] 1
radar_8km RainViewer 192 8 1536 [0] 1
satellite_8km EUMETSAT 192 8 1536 [-120,-105,…,-60] 11
gfs_16km GFS 96 16 1536 [0] 122
gfs_forecast_16km GFS 96 16 1536 [60,120,…,480] 1
xyz_4km DEM 256 4 1024 N/A 3
minute_4km - 256 4 1024 N/A 1
[…]
A.3 Split Generation
Data spans from December 2023 to November 2024 and is partitioned into training, validation, and testing sets using multi-day cycles (12, 2, and 2 days, respectively). A 12-hour blackout period between cycles prevents data leakage.
Samples are generated using a sliding window along the temporal dimension for non-overlapping $512\times 512\mathrm{km}^{2}$ patches, ensuring alignment with the timesteps and spatial dimensions in Table 3. The training set includes 1,063,658 samples, each with at least 50% radar coverage. During training, we apply uniform random offsets of ±256km horizontally and vertically to patches to make the training dataset as large as possible. Validation and test sets comprise 5,120 and 15,360 randomly selected samples, respectively, including challenging cases with radar coverage below 50%. Although evaluation is limited to pixels with radar coverage, boundary regions remain more challenging to predict due to the absence of nearby radar coverage.
Appendix B Metrics
B.1 Metric Definitions
Critical Success Index (CSI)
CSI, also known as the Threat Score, evaluates the accuracy of event detection by comparing the correctly predicted precipitation events to all events that were either predicted or actually occurred. It balances misses and false alarms, making it particularly useful in rare-event forecasting like heavy rainfall. A value of 1 indicates perfect accuracy, while 0 indicates no skill.
$\text{CSI}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},$ (8)
where TP, FP, and FN are true positives, false positives, and false negatives, based on the threshold (rate $\geq$ threshold). The thresholds at which the metric is evaluated are 0.5, 1, 2, 5, and 10 mm/h.
Continuous Ranked Probability Score (CRPS)
CRPS measures the accuracy of the entire predicted cumulative distribution function (CDF) relative to the observed outcome. It generalizes MAE for probabilistic forecasts, with lower values indicating better forecast reliability, sharpness, and calibration. It rewards predictions that assign high probability to the correct intensity range.
$\text{CRPS}=\sum_{c=1}^{|I|}[(P(R_{t}<\min(I_{c}))-\mathds{1}(R_{t}<\min(I_{c}))]^{2}\times|I_{c}|,$ (9)
where $c$ iterates over all intensity classes, $\min(I_{c})$ is the lower end of class $c$, $R_{t}$ is the target radar, and $P(R_{t}<\min(I_{c}))=1-P(R_{t}\geq\min(I_{c})$, based on the model probability $P_{t,c}$.
Fraction Skill Score (FSS)
FSS provides an estimate of spatial accuracy by comparing the forecast and observed precipitation fields over a local neighborhood, rather than point-wise accuracy. FSS is particularly useful for evaluating high-resolution forecasts, as it accounts for small spatial displacements in predicted precipitation. A perfect forecast yields an FSS of 1, while lower values indicate poorer spatial agreement.
$\text{FSS}=1-\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}(F_{i,j}-O_{i,j})^{2}}{\sum_{i=1}^{H}\sum_{j=1}^{W}F_{i,j}^{2}+\sum_{i=1}^{H}\sum_{j=1}^{W}O_{i,j}^{2}},$ (10)
where $F_{i,j}$ and $O_{i,j}$ refer to the fraction of predicted positives and fraction of observed positives, respectively, in the neighborhood of the $(i,j)$ pixel. This metric is computed at different thresholds (0.5, 1, 2, 5, and 10 mm/h) and neighborhood sizes (2km, 10km, and 20km). This metric also requires binarization based on the computed thresholds.
Frequency Bias Index (FBI)
FBI quantifies whether a model tends to overforecast or underforecast precipitation events. A value greater than 1 indicates overprediction, while a value less than 1 indicates underprediction. While a bias of 1 is ideal, it doesn’t imply the forecast is accurate—just that the frequency of forecasted events matches the observed frequency.
$\text{FBI}=\frac{\text{TP}+\text{FP}}{\text{TP}+\text{FN}},$ (11)
where TP, FP, and FN are true positives, false positives, and false negatives, based on the threshold (rate $\geq$ threshold). The thresholds at which the metric is evaluated are 0.5, 1, 2, 5, and 10 mm/h.
B.2 Additional Evaluation
Table 4: Critical Success Index (CSI) across different thresholds for precipitation forecasting models. Our model RainPro-8 outperforms all other models across different rain intensities.
CSI ($\uparrow$)
Model 0.2 1.0 2.0 5.0 10.0
ours 0.447 0.344 0.296 0.204 0.146
metnet-3* 0.437 0.335 0.287 0.194 0.136
gfs 0.253 0.161 0.100 0.027 0.009
pysteps 0.271 0.180 0.141 0.090 0.062
nowcastnet 0.261 0.118 0.082 0.036 0.018
earthformer 0.265 0.125 0.088 0.039 0.015
simvp 0.256 0.154 0.116 0.047 0.023
Table 5: Fraction Skill Score (FSS) across different thresholds for precipitation forecasting models with a neighborhood size of 10km. Our model RainPro-8 outperforms all other models across different rain intensities.
FSS-10km ($\uparrow$)
Model 0.2 1.0 2.0 5.0 10.0
ours 0.728 0.639 0.590 0.469 0.363
metnet-3* 0.716 0.625 0.572 0.445 0.336
gfs 0.489 0.370 0.268 0.101 0.041
pysteps 0.552 0.430 0.369 0.283 0.225
nowcastnet 0.492 0.256 0.191 0.102 0.057
earthformer 0.509 0.279 0.211 0.116 0.056
simvp 0.485 0.338 0.277 0.134 0.075
Table 6: Frequency Bias Index (FBI) across different thresholds for precipitation forecasting models. Our model, RainPro-8, has a tendency to over-predict high-precipitation events, in contrast to GFS, NowcastNet, Earthformer, and SimVP, which consistently under-predict rainfall. PySTEPS appears the most balanced in terms of bias, with over- and under-predictions occurring at similar rates—though this does not translate to higher forecast accuracy.
FBI ($\approx 1$)
Model 0.2 1.0 2.0 5.0 10.0
ours 1.140 1.137 1.142 1.268 1.636
metnet-3* 1.124 1.136 1.160 1.364 1.821
gfs 1.602 1.177 0.700 0.292 0.127
pysteps 0.899 0.995 0.914 1.029 1.078
nowcastnet 0.398 0.165 0.116 0.052 0.026
earthformer 0.397 0.182 0.135 0.063 0.027
simvp 0.384 0.244 0.197 0.074 0.034
The following plots expand on the previous tables in terms of lead times.
Figure 5: Critical Success Index (CSI) for Europe-wide precipitation forecasting models across different thresholds and lead times.
Figure 6: Fraction Skill Score (FSS) for Europe-wide precipitation forecasting models across different thresholds and lead times with neighborhood size of 10km.
Figure 7: Frequency Bias Index (FBI) for Europe-wide precipitation forecasting models across different thresholds and lead times.
Appendix C Input Attribution with Integrated Gradients
The following plots show the results obtained on the input attribution by using Integrated Gradients from which the key findings were obtained in Section 4.4. Integrated Gradients estimates feature importance by interpolating between a baseline and the actual input, averaging model gradients along this path. The final attribution is obtained by scaling these averaged gradients by the input difference, quantifying each feature’s contribution to predictions. To ensure meaningful attribution, we use the minimum value of each feature as the baseline, as this results in near-zero precipitation probabilities. We then aggregate feature attributions across space, time, and samples, yielding a single importance score per feature.
Figure 8: Attribution of radar_4km input timesteps over forecast lead time.
Figure 9: Attribution of radar_8km over forecast lead time.
Figure 10: Attribution of satellite_8km input timesteps over forecast lead time.
Figure 11: Attribution of satellite_8km input variables over forecast lead time.
Figure 12: Attribution of gfs_16km input variables over forecast lead time.
Figure 13: Attribution of gfs_forecast_16km input timesteps over forecast lead time.
Appendix D Visualizations
This section displays a Europe-wide comparison of forecast models and probability maps for RainPro-8 at various precipitation levels, based on three different origins.
Figure 14: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-01-23 11:20 UTC.
Figure 15: Probability map with corresponding ground truth for four different rain intensities and two different lead times with origin at 2024-01-23 11:20 UTC.
Figure 16: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-04-19 06:00 UTC.
Figure 17: Probability map with corresponding ground truth for four different rain intensities and two different lead times with origin at 2024-04-19 06:00 UTC.
Figure 18: Sample ground truth, RainPro-8, MetNet-3*, GFS, and PySTEPS, NowcastNet, Earthformer, and SimVP forecasts at selected lead times with origin at 2024-06-28 11:20 UTC.
Figure 19: Probability map with corresponding ground truth for four different rain intensities and two different lead times with origin at 2024-06-28 11:20 UTC.
Appendix E Intensity Classes
Table 7 presents the intensity classes used by our model along with their distribution across different splits. Note that both the validation and test sets contain significantly more missing values due to patches with radar coverage below 50% to obtain Europe-wide forecasts.
Table 7: Intensity classes and distribution for each split.
Bucket (mm/h) Train (%) Validation (%) Test (%)
$[0.0,0.1)$ 79.64 57.44 57.05
$[0.1,0.2)$ 1.44 1.00 1.10
$[0.2,0.4)$ 2.25 1.62 1.87
$[0.4,0.6)$ 0.84 0.62 0.74
$[0.6,0.8)$ 0.51 0.37 0.44
$[0.8,1.0)$ 0.50 0.36 0.44
$[1.0,2.0)$ 0.84 0.63 0.73
$[2.0,3.0)$ 0.38 0.30 0.35
$[3.0,4.0)$ 0.19 0.15 0.17
$[4.0,5.0)$ 0.17 0.13 0.14
$[5.0,6.0)$ 0.06 0.05 0.05
$[6.0,7.0)$ 0.05 0.04 0.04
$[7.0,8.0)$ 0.04 0.04 0.03
$[8.0,9.0)$ 0.03 0.02 0.02
$[9.0,10.0)$ 0.02 0.02 0.02
$[10.0,15.0)$ 0.03 0.03 0.02
$[15.0,20.0)$ 0.02 0.02 0.01
$[20.0,25.0)$ 0.01 0.01 0.01
$\geq 25.0$ 0.02 0.02 0.01
missing 12.97 37.15 36.78
Appendix F GFS Input Variables
Table 8: Selected GFS variables as input to the model with their corresponding levels.
Var. Description Levels
4LFTX Best (4 layer) Lifted Index surface
ABSV Absolute Vorticity 100mb, 250mb, 500mb, 850mb, 1000mb
CAPE Convective Available Potential Energy surface, 180-0mb, 90-0mb, 255-0mb
CFRZR Categorical Freezing Rain surface
CICEP Categorical Ice Pellets surface
CIN Convective Inhibition surface, 180-0mb, 90-0mb, 255-0mb
CLMR Cloud Mixing Ratio 250mb, 500mb, 850mb, 1000mb
CPOFP Percent frozen precipitation surface
CRAIN Categorical Rain surface
CSNOW Categorical Snow surface
CWAT Cloud Water entire atmosphere
DPT Dew Point Temperature 2m
DZDT Vertical Velocity (Geometric) 100mb, 250mb, 500mb, 850mb, 1000mb
GRLE Graupel 100mb, 250mb, 500mb, 850mb, 1000mb
GUST Wind Speed (Gust) surface
HGT Geopotential Height 100mb, 250mb, 500mb, 850mb, 1000mb, surface, trop.
HPBL Planetary Boundary Layer Height surface
ICEG Ice Growth Rate 10m
ICETK Ice Thickness surface
ICMR Ice Water Mixing Ratio 250mb, 500mb, 850mb, 1000mb
LCDC Low Cloud Cover low cloud layer
LFTX Surface Lifted Index surface
MCDC Medium Cloud Cover middle cloud layer
MSLET (Eta model reduction) mean sea level
PLPL Pressure of level from which parcel was lifted 255-0mb
PRATE Precipitation Rate surface
PRES Pressure surface, trop.
PRMSL Pressure Reduced to MSL mean sea level
PWAT Precipitable Water entire atmosphere
REFC Composite reflectivity entire atmosphere
RH Relative Humidity 100mb, 250mb, 500mb, 850mb, 1000mb, 2m, entire atmosphere
RWMR Rain Mixing Ratio 100mb, 250mb, 500mb, 850mb, 1000mb
SNMR Snow Mixing Ratio 100mb, 250mb, 500mb, 850mb, 1000mb
SNOD Snow Depth surface
SPFH Specific Humidity 100mb, 250mb, 500mb, 850mb, 1000mb, 2m
TCDC Total Cloud Cover 250mb, 500mb, 850mb, 1000mb, entire atmosphere
TMP Temperature 100mb, 250mb, 500mb, 850mb, 1000mb, surface, 2m, trop.
UGRD U-Component of Wind 100mb, 250mb, 500mb, 850mb, 1000mb, 10m, trop.
USTM U-Component Storm Motion 6000-0m
VGRD V-Component of Wind 100mb, 250mb, 500mb, 850mb, 1000mb, 10m, trop.
VIS Visibility surface
VSTM V-Component Storm Motion 6000-0m
VVEL Vertical Velocity (Pressure) 100mb, 250mb, 500mb, 850mb, 1000mb
VWSH Vertical Speed Shear trop.
WEASD Water Equivalent of Accumulated Snow Depth surface
Appendix G Compute Resources
Training is performed on an NVIDIA H100 80GB SXM5 GPU with 26 vCPUs and 225 GiB RAM. Table 9 shows the training time for each experiment including baselines and ablation studies.
Table 9: .
Experiment Training Time (h)
RainPro-8 13:36
MetNet-3* 18:16
NowcastNet 6:48
Earthformer 24:03
SimVP 6:36
RainPro-8_ce 15:52
RainPro-8_clt 18:06
RainPro-8_noltw 13:21
RainPro-8_nomask 13:20
RainPro-8_sources 13:22